OpenStudy (anonymous):

solve over domain -3pi/2 greater or equal to x < pi/2 for (cos2x)(cosx) - (sin2x)(sinx) = 0 trig identity: cos(alpha + beta)

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

cos 3x = 0

OpenStudy (anonymous):

@soapia yes u are correct Cos3x = 0...... But can u write the domain again???

OpenStudy (anonymous):

-3pi/2 greater or equal to x < pi/2

OpenStudy (anonymous):

u want 2 say \[-3\pi /2\le x <\pi /2\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

See general solution for \[\cos x = 0\] is \[x = 2n \pi \pm \pi / 2\] So if we check the domain we get \[3x = -\pi / 2, -\pi / 2, \pi / 2\] etc. So, \[x= -\pi / 2, -\pi / 6, \pi / 6\]

OpenStudy (anonymous):

I think the general solution is: \[\cos (x ) = 0\] \[x = (2n + 1) \frac{\pi}{2} \qquad where \; \; n \in \; \mathcal{Z}\]

OpenStudy (anonymous):

Or as @lordcyborg says: it should be: \[x = n \pi + \frac{\pi}{2}\]